How much NAOH is required to prepare 50 mL of aqueous solution with 70 mg of Na+ ions per mL of the solution?
1
Expert's answer
2019-05-23T07:58:16-0400

50 mL aqueous solution with concentration of Na+ ions 70mg/mL will contain:

"m(Na^+)=c(Na^+) \cdot V(Na^+) = 70\frac{mg}{mL} \cdot 50 mL = 3500 mg = 3.5 g"

The amount of Na+ ions in moles:

"n(Na^+) = \frac{m(Na^+) }{M(Na^+) } = \frac{3.5g}{23\frac{g}{mol}} = 0.1522 mol"

The same amount of NaOH is required to prepare solution:

"n(NaOH) = n(Na^+)=0.1522 mol"

Mass of NaOH is:

"m(NaOH) = n(NaOH) \cdot M(NaOH) = 0.1522 mol \cdot 40 \frac{g}{mol} =6.09 g"